GreenHab Report – April 29th

Crew 211 GreenHab Report – 29-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 2

Shade cloth: 80%

Average temperatures:

Low: 16°C

High: 37°C

Hours of supplemental light: 0

Daily water usage for crops: 5 gallons

Water in Blue Tank: 80 gallons

Time(s) of watering for crops: 14:00

Changes to crops: All plants look healthy. Nevertheless the tomato plant of the GreenHab still seems a little stressed (curved leaf), I will water it accordingly over the next couple of days. There are more flowers on the Parsley plants, Lettuce (old seed) and even on radish. Cucumber plants grow well.

Narrative: It was cooler today, the rain has helped to decrease the temperature. There was a visit of the PBS in the GreenHab.

Harvest: 6.3g Parsley, 9.8g Basil

Support/supplies needed: I won’t have enough water for 1 week long, as it was really hot this past week, and I had to water plants accordingly.

Science Report – April 29th

Science Report 29 April 2019
Crew 211 – UCL to Mars
Submitted by Science Officer Maxime Bernard

Commander Carl Dahlqvist

All the Hardware is done and tested and all the CubeSat are all mounted.
Only the software still need to be tested.

XO Simon Collignon

Arrive à dessiner en temps réel le tracé d’objets (RTLS: Real Time Location System)

HSO Benjamin Flasse

This morning Benjamin gathered the data from the sleep monitoring he did on one of the crew member.

Right after that, as every morning, he measured everybody’s weight, tension and brain awakeness and all the usual mensurations.
All CFFF, blood pressure and heart rate graph are done.

Crew Astronomer Eleonore Lieffrig

Tried to deal with the Musk observatory and the tracking system
Finished to compute the Brownian motion generator and the Peclet Number.

Crew Engineer Julien Amalaberque

Managed to align two video flux (color and depth) to successfully compute the real time camera position.
Next step will be to find a way to fix it to the drone.

Crew Biologist Chloé Peduzzi

Every spirulina cultures are going to be checked using an axenicity test which will last 3 days and will allow her to know if any contamination happened.

Crew Journalist Nathan Pechon

We had a meeting during which we discussed about the symbol on Mars(flag,..), the religion and thought freedom and finally about the right for informations.

Crew Scientist Maxime Bernard

Still trying to figure out where the statistic counting problem could come from.
Only few muons can be seen.

EVA Report – April 29th

EVA report #8

Date: 29 April 2019 from 9:00 to 12:00 am. But as the PBS crew was there, it was delayed from 10:h30 to 12:30 am.

The purpose of EVA was Mars exploration and collection of soil samples in order to grow plants in a Martian environment for the biologist (Chloé Peduzzi) experiment, but also to show to the PBS team a glimpse of a typical EVA.

Destination: Latitude: 4248000, longitude: 519000 (UTM NAD27 CONUS)

EVA Participants: Chloé Peduzzi (commander of the EVA), Benjamin Flasse (medical officer), Eléonore Lieffrig, Nathan Pechon, Lys Adeola (PBS host)

Roads and routes per MDRS Map: Cow Dung Road 0110 south direction for about 1.5 miles.

Mode of travel: Driving and hiking with Spirit and Curiosity (and PBS’s car).

Narrative : After an immersive suit of for Lys, we headed off towards our destination with Spirit, Curiosity and the PBS’s car and drone following us. We managed to take a walk in the planned area, and the EVA commander Chloé Pedduzi harvested some soil samples. After climbing a small hill, we discovered a beautiful Mars-like view. Both the PBS team and us were amazed!

We all hiked in this beautiful landscape and found what seems to be a cavity, we thought that we had just found a shelter ! But after a better and closer looked, we discovered that it was just a small and not big enough cavity.

Journalist Report – April 29th

Today was our eighth simulation day. We woke up at 7:30 am and we had breakfast at 8:00. At 9 am, we received KUED Channel 7. PBS, a local tv interviewed us. Carl welcomed the PBS team in spacesuit outside.

Liz, a TV presenter, was very interested by the station. We showed all rooms of the Hab, the greenhab, the Ram, science dome and the solar telescope. She interviewed us about our martian life. After that, she went with us in EVA. She wore a space suit and become an astronaut during one hour. We found the entrance of what seemed to be a grand cavity which can be used as a shelter in case of strong solar winds.

After the EVA, we ate with the TV team. After meal, we discussed about Nathan’s experience and his Martian constitution. We debated about religions and thought freedom on Mars.

At 2 pm Liz interviewed Carl and Eleonore. She asked why science is important, is it difficult to live in the MDRS, … ?

In the same time we worked on our experiments. Read the scientist report for more informations.

At 6 pm, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. After the Cap Com, we will eat leftovers for dinner.

Like usual, we will do a crew meeting to speak about the day.

Operations Report – April 29th

Crew 211 Operations Report 29-APR-2019

SOL: 8

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 13:00

On (last night): 6:30 pm

Off (this morning): 7:30 am

Solar— SOC 71 % at 14:30 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~290 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 80 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 15 gallons

Toilet tank emptied: yes

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 90.2h

Beginning charge: 100

Ending charge: 55

Currently charging: yes

Opportunity rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours: 92.2h

Beginning charge: 100

Ending charge: 72

Currently charging: yes

Notes on rovers: Nothing to report

Reason for use: EVA with PBS

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some basil and parseleaf was harvested

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support:

* Some additional water for the GreenHab would be useful.

Sol Summary – April 29th

Crew 211 Sol 8 Summary Report 29-APR-2019

Sol: 8

Summary Title: Mars on TV

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After a quick breakfast, the commander welcomed the PBS team in spacesuit outside of the Hab. We then made a tour of the MDRS campus with the TV crew, each member of the UCL to Mars team presenting part of the infrastructures and discussing his/her experiment. Half of the crew left then for an EVA with Liz Adeola to Cow Dung Road between the Robert’s Rock Garden and Zurbin’s head. The EVA team got back at noon and we all had a nice lunch together. The crew commander and the crew astronomer were then interviewed for about one hour. All the crew members went then back to work once the TV crew left the campus.

Look Ahead Plan: Depending on the weather, we plan to start the day with an EVA to Kissing Camel. We will then work on our respective experiments for the rest of the day preparing the first experiment related EVA for the commander and vice-commander which should be organized on Wednesday.

Anomalies in work: No issues.

Weather: Showers, mostly cloudy and windy

Crew Physical Status: Good

EVA: Half of the crew explored the west part of Cow Dung Road with the PBS crew. Liz Adeola was part of the EVA team, as we her a spacesuit for the occasion. The EVA lasted for about an hour. The PBS team did a few drone footages and interviewed the EVA team members.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site